OpenStudy (anonymous):

Factorize : 3t^3 +4t^2 - 5t - 2 = 0

OpenStudy (agent47):

What does factorize mean?

OpenStudy (anonymous):

search on google.

OpenStudy (agent47):

Why can't you search on google?

OpenStudy (anonymous):

b'coz i know what it means.

OpenStudy (agent47):

Well why can't you answer the question then? I'm just asking for additional info.

OpenStudy (anonymous):

factorize means to express an equation as a product of it factors.

OpenStudy (agent47):

So factorize means factor? 3x^3 +4t^2 - 5t - 2 = 0

OpenStudy (anonymous):

-.-,

OpenStudy (anonymous):

@Preetha Help!!

OpenStudy (agent47):

You can't factor that equation btw.

myininaya (myininaya):

Do you know synthetic division?

myininaya (myininaya):

Do you know how to spot possible rational zeros?

OpenStudy (anonymous):

is that supposed to be an x or a t at the beginning?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but i know synthetic division

OpenStudy (anonymous):

steps pls.

myininaya (myininaya):

Hey @robtobey he needs to understand how to get those factors

myininaya (myininaya):

Also @shubham.bagrecha which equation are we solving?

OpenStudy (anonymous):

both are same.

OpenStudy (anonymous):

anyone of your choice

myininaya (myininaya):

They aren't the same ...

OpenStudy (anonymous):

the 1st eq. is obtained by movin 1/t to LHS and taking LCM.

myininaya (myininaya):

\[3t^2+4t-5=\frac{1}{t}\] Assume t does not equal 0 Multiply t on both sides \[3t^3+4t^2-5t=1\] Now subtract 1 on both sides

OpenStudy (anonymous):

where is that robot guy?? he deleted his post.

myininaya (myininaya):

@shubham.bagrecha yes that is how you multiply both sides by t

myininaya (myininaya):

now what is 1/t * t

myininaya (myininaya):

not 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but still what are the factors

OpenStudy (anonymous):

????

myininaya (myininaya):

The only way I see is approximating the solutions Like you go through the possible rational zeros: \[\pm1 , \pm 3 , \pm \frac{1}{3}\] you will see that none of these work

OpenStudy (anonymous):

Do you know log operations?

OpenStudy (anonymous):

@myininaya the 2 in the equation is correct.

myininaya (myininaya):

which equation are we looking at?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

i wrote the second one wrong.

myininaya (myininaya):

Ok. So do you know how to spot the rational zeros now?

OpenStudy (anonymous):

\[t(3t^{2}+4t-5)=2\] So can we write it as-

OpenStudy (anonymous):

\[t/2(3t^{2}+4t-5=0)\]

myininaya (myininaya):

Like you look for possible factors of -2 and 3 And then do the possible factors of -2 over 3 Like this : \[\pm 1, \pm 2 , \pm 3 ,\pm \frac{2}{3}\] Whenever you have \[ax^n+...+b=0\] To find possible rational zeros you do you write the poss. factors of a , of b , and then poss of b / poss of a So We can start we any of those possible zeros and use synthetic division to find the other quadratic factor if a rational zero exist

myininaya (myininaya):

You can't do that @shubham.bagrecha Because 2/2=1 not 0

OpenStudy (anonymous):

ok

myininaya (myininaya):

So you said you knew how to do synthetic division right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Try the possible rational zeros I have above and see if one works

myininaya (myininaya):

one of them* one not to be confused with that 1 I have up there

OpenStudy (anonymous):

ok

myininaya (myininaya):

oops missed a possible rational zero \[\pm \frac{1}{3}\]

myininaya (myininaya):

some*

myininaya (myininaya):

\[poss. factors of -2, poss. factors of 3, \frac{ poss. factors of -2}{poss. factors of 3}\]

myininaya (myininaya):

just explaining why those were missing from above

myininaya (myininaya):

Have you tried 1 yet ?

OpenStudy (anonymous):

no

myininaya (myininaya):

What did you get when you tried 1?

myininaya (myininaya):

@shubham.bagrecha you are almost there :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

leaving it here.sorry

myininaya (myininaya):

you are almost there i promise

myininaya (myininaya):

Like doing t=1 1 | 3 4 -5 -2 | -------------------------- | Bring down the 3 1 | 3 4 -5 -2 | -------------------------- 3 | Do that 1 times that 3 1 | 3 4 -5 -2 | 3 -------------------------- 3 | Add the ones in the column 4+3 1 | 3 4 -5 -2 | 3 -------------------------- 3 7 | Can you finish this part?